OpenStudy (anonymous):

HELP!! Simplify log8 7 – log8 s + log8 t – log8 4 HELP!! Simplify log8 7 – log8 s + log8 t – log8 4 @Mathematics

OpenStudy (anonymous):

log base 8?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

\[\log(\frac{7t}{4s})\]

OpenStudy (anonymous):

are u sure?

OpenStudy (anonymous):

ones with plus go in the numerator, ones with minus go in the denominator

OpenStudy (anonymous):

oh ya

OpenStudy (anonymous):

Ok so its like putting the first half into a fraction and the second part another fraction and multiply across?

OpenStudy (anonymous):

yeah i am pretty sure... of course it is true that \[\log_8(4)=\frac{2}{3}\] but i doubt this is part of the problem

OpenStudy (anonymous):

you forgot to add the base 8 in ur first answer

OpenStudy (anonymous):

\[\log(a)+\log(b)=\log(ab)\] and \[log(a)-\log(b)=\log(\frac{a}{b})\]

OpenStudy (anonymous):

base is not relevant to this problem

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

So would this: Solve log7 (x – 8) – log7 3 = log7 21 be x = 71

OpenStudy (anonymous):

x is 71?

OpenStudy (anonymous):

I don't think so

OpenStudy (anonymous):

(x-8)/3=3

OpenStudy (anonymous):

x-8=9

OpenStudy (anonymous):

you need an x in your answer. it would be \[\log(\frac{x-8}{3})\]

OpenStudy (anonymous):

x=17

OpenStudy (anonymous):

Oh so am I wrong?

OpenStudy (anonymous):

\[\frac{x-8}{3}=21\] \[x-8=63\] \[x=71\]

OpenStudy (anonymous):

But how did you get (x - 8) / 3 = 3 its suppose to equal 21?

OpenStudy (anonymous):

I think I am wrong

OpenStudy (anonymous):

yes it is equal to 21

OpenStudy (anonymous):

71 is correct

OpenStudy (anonymous):

I have another question, but I will put it on the side so i can give you guys another medal if you help me ^^

OpenStudy (anonymous):

Ya i messed up

OpenStudy (anonymous):

oh who cares abt the medals???

OpenStudy (anonymous):

I just feel bad if I dont give you guys what you deserve for taking ur time to help me haha